# A noninvasive and nonadiabatic quantum Maxwell demon

### v1

@@ Abstract
Introduction—After decades of resting as a conceptual problem understood in terms of information erasure [1, 2], the interest on the Maxwell demon, a mechanism able to decrease the entropy of a system without acting directly on its particles [3], has been revived in the last years. The way has been paved by the development of nanoscale technologies, with their access to microscopic degrees of freedom, bringing proposals and experimental realizations where the action of Maxwell demon devices [4, 5, 6, 7, 8] can actually be measured. This is so that a rich demonology has come out in the last years based on different platforms (e.g., quantum dots, ions, qubits), classified depending on whether they need active measurement and feedback on the system [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] or they work autonomously by means of mutual information flows in nonequilibrium configurations [19, 20, 21, 22, 23]. Information heat engines have also been proposed [24, 25, 26, 27, 28] and realized [29, 30] which are based on the exchange of mutual information [31] but cannot avoid the flow of heat.
@@
FIG. 1: Three steps of the demon operation inducing transport through a DQD coupled to reservoirs L and R: detection (of whether there is a particle in the DQD), operation (perform a voltage pulse that exchanges the energies of the left and right dot levels), and reset. The demon performs a continuous measurement of the total charge, NN and performs feedback operations when NN changes.
@@
In the quantum regime, qubit based proposals [32, 33, 34, 35, 36, 37, 38, 39, 40] and proof of concept experiments [41, 42, 43, 44, 45] use quantum measurements [46, 47] and unitary gates. They involve either local projective measurements (so decoherence and work injection are unavoidable) or that the measurement and the feedback rotation are performed in different bases of the same qubit (so they cannot coexist). Here we propose a Maxwell demon that exploits the quantum nature of the two demonic actions (detection and feedback) combining noninvasive continuous measurements and efficient driven tunneling [48] in a transport setting. It augments the qubit Hilbert space, ${\bf 1}=\{|l\rangle,|r\rangle\}$, by an empty state $|0\rangle$ that couples it to the work (or heat) extraction reservoirs and allows for continuous detection without perturbing the qubit.
@@
We exemplify this with a double quantum dot (DQD) charge qubit [49, 50] with interdot coupling $\lambda$ mediating the particle flow between two reservoirs (L and R) and coupled to a charge detector (e.g., a quantum point contact [51, 52, 53, 54], a single electron transistor [55] or a resonator [56, 57, 58, 59, 60]), see Fig. 1. The two states in $\bf 1$} are initially energetically split: $\varepsilon_{l}=\varepsilon_{0}$ and $\varepsilon_{r}=\varepsilon_{0}+\Delta_{\varepsilon}$, so the occupation of $\lvert l\rangle$ is favored. Upon the detection of an electron (most likely coming from L) the demon exchanges the level energies in a time $\tau_{d}$. During the level crossing, the electron tunnels to the other quantum dot with a Landau-Zener-Stückelberg-Majorana (LZSM) probability $P_{LZ}$ that is controlled by the driving speed, $\Delta_{\varepsilon}/\tau_{d}$ [61, 62, 63, 64]. Swapping both the energy and the charge of the quantum dots, the successful operation hence conserves energy and enables transport: the electron is now able to tunnel out to R. Finally the initial state is reset. Provided $\Delta_{\varepsilon}\gg\lambda$, efficient $|l\rangle\rightarrow|r\rangle$ transfer is possible in the adiabatic ($\tau_{d}\gg\hbar/\lambda$) regime [65] as well as for particular speeds [66, 67]. Fixing $\tau_{d}$ also makes the system more autonomous and less detector-dependent. LZSM tunneling is routinely used in DQDs to enable quantum gates [68, 69, 70, 71, 72, 73, 74, 75], and to probe coherence [76, 77, 78, 79, 80, 81, 82], also in qubits [83, 84, 85, 86]. Note that charge pumps generate transport using periodic drivings [87, 88, 89, 90] without measuring, or conditioned feedback [91], in both cases with the drive performing work on the system [92]. To avoid this, the information of when the DQD gets occupied is essential.
@@
The challenge is to measure continuous and noninvasively in a way that the state coherence is not disturbed during the driving operation. Protocols conditioned on resolved knowledge of which dot ($l$ or $r$) is occupied [16] impede the coherent driving and are affected by heat injection [93, 94, 95]. We avoid these drawbacks by imposing that the demon has only a limited access to the state of the system: it detects whether the DQD is empty ($N=0$) or occupied ($N=1$), but not in which quantum dot the electron is. Hence, using the detection operator $\hat{N}_{1}=\lvert l\rangle\langle l\rvert{+}\lvert r\rangle\langle r\rvert$ (the detector is equally coupled to both dots), the coherences between states $|l\rangle$ and $|r\rangle$ are respected during the operation step, even if the detector continues measuring. Note that the total charge $N=\langle\hat{N}\rangle$ plays the role of the memory of the demon in a Szilard engine [96, *szilard_on_1964]. By measuring $N=1$, the demon will hence not be certain that the electron is in the left quantum dot, as desired, but can prepare the system such that this is most likely the case if $\varepsilon_{0}\ll\varepsilon_{0}+\Delta_{\varepsilon}$.
@@
There is a price for indeterminacy: if the driving goes wrong (with probability $1-P_{LZ}$) the demon performs work and heats the system, which limits its performance far from the adiabatic regime; if the electrons tunnels from R, the demon transports it in the wrong direction.
@@
$\hat{H}_{\mathrm{0}}(t)=\sum_{\alpha=l,r}\varepsilon_{\alpha}(t)\hat{n}_{\alpha}+\lambda\left(\hat{d}_{\mathrm{l}}^{\dagger}\hat{d}_{\mathrm{r}}+\hat{d}_{\mathrm{r}}^{\dagger}\hat{d}_{\mathrm{l}}\right).$ (1)
The time dependence is due to the demon action, which D according to the following protocol (see Fig. 1): (i) if $N=0$, $\varepsilon_{l}=\varepsilon_{0}$ and $\varepsilon_{r}=\varepsilon_{0}+\Delta_{\varepsilon}$ (detection I stage); (ii) by detecting a change to $N=1$ at $t=t_{k}$, the demon swaps the energies of the dots via gate voltages in a time $\tau_{d}$: $\varepsilon_{l}(t)=\varepsilon_{0}+\Delta_{\epsilon}(t-t_{k})/\tau_{s}$, $\varepsilon_{r}(t)=\varepsilon_{0}+\Delta_{\epsilon}(\tau_{d}t-t_{k})/\tau_{d}$ (driving stage); (iii) at $t=t_{k}+\tau_{d}$, $\varepsilon_{l}=\varepsilon_{0}+\Delta_{\varepsilon}$ and $\varepsilon_{r}=\varepsilon_{0}$ is maintained until $N=0$ is detected (detection II stage); (iv) finally a fast drive (compared to $1/\Gamma$) is performed to restore the initial state (reset stage). If $\varepsilon_{0}<\mu_{L}$ and $\mu_{R},k_{\text{B}}T\ll\varepsilon_{0}+\Delta_{\varepsilon}$, where $\mu_{j}$ is the electrochemical potential of reservoir $j=L,R$, tunneling into the DQD from R will be exponentially suppressed by a rate $\propto e^{-(\varepsilon_{0}+\Delta_{\varepsilon}-\mu_{R})/k_{\text{B}}T}$, so the demon can assume that most electrons come from L. To avoid further sources of errors, the demon also decouples the system from the reservoirs during the driving step, which can be done with additional gate voltages. We parametrize this as $\Gamma(t)=\Gamma\chi(t)$, where $\chi=0$ during the driving stage and 1 otherwise. In a successful sequence, the electron tunnels from L to $l$ at $t=t_{k}$, is transferred to $r$ at $t=t_{k}+\tau_{d}$ and then tunnels out to R. In that case, a particle is transferred across the system, heat $\mu_{j}-\varepsilon_{0}$ is extracted from both reservoirs and the demon has performed no work on the system. All the thermodynamic flows can be reversed in the system. The resource is in this case the amount of information generated in the detection apparatus to maintain the continuous measurement [98, 99, 100, 101].
We assume a fast and errorless detector. The evolution of the conditioned reduced density matrix of the DQD, $\rho_{\gamma}$ [102], for a finite time trajectory based on a sequence of measurements records ($\gamma_{(0,\tau)}$) is given by the stochastic master equation [47]:
@@
where $L_{j}^{+}=\sqrt{\Gamma f_{j}}\lvert\alpha_{j}\rangle\langle 0\rvert$ and $L_{j}^{-}=\sqrt{\Gamma(1{-}f_{j})}\lvert 0\rangle\langle\alpha_{j}\rvert$ are jump operators associated with the tunneling in ($i=+$) and out ($i=$-) of the DQD via reservoir $j$, and $f_{j}=[e^{(\varepsilon_{\alpha_{j}}-\mu_{j})/k_{\text{B}}T}+1]^{-1}$ is the Fermi-Dirac distribution ($\alpha_{L}=l$, $\alpha_{R}=r$), which takes two values: $f_{j}^{0}$ and $f_{j}^{\Delta}$ depending on whether $\varepsilon_{\alpha_{j}}=\varepsilon_{0}$ or $\varepsilon_{0}+\Delta_{\varepsilon}$. Following Eq. (4), an initially empty DQD is updated to a state $\rho_{\gamma}^{\prime}=(f_{L}\lvert l\rangle\langle l\rvert+f_{R}\lvert r\rangle\langle r\rvert)/(f_{L}+f_{R})$. In the case $\Delta_{\varepsilon}\gg\mu_{R}+k_{\text{B}}T$ and $\mu_{L}-\varepsilon_{0}\sim k_{\text{B}}T$, the ratio $f_{L}/f_{R}\gg 1$ guarantees that the protocol will start with a state $\rho_{\gamma}^{\prime}\approx\lvert L\rangle\langle L\rvert$, with exponentially suppressed deviations, and $1-f_{j}$ is finite to enable the electron tunneling out to R after the driving. Additionally, the condition $\Delta_{\varepsilon}\ll\varepsilon_{0}$ allows our local description to be accurate [103].
FIG. 2: (a) Average particle current through the DQD as a function of the applied electrochemical potential bias and the speed of the energy ramp, with cuts along the marked dotted lines plotted in the lateral panels. $\Gamma=\Omega/\hbar$, $\lambda=0.5\Omega$, $T_{l}=T_{r}=0.75\Omega/k_{\rm B}$, $\varepsilon_{0}=-0.5\Omega$, $\mu_{r}=0$, $\Delta=6\Omega$. The dashed line in the $\Delta\mu=0$ panel plots $\tau_{d}+1/\Gamma f^{0}(1-f^{0})]^{-1}$. (b) Histograms of the time-averaged current $I_{\gamma}\equiv\langle I^{\gamma}(t)\rangle_{t}$ of $10^{6}$ trajectory computed during a time $t_{run}=10^{5}\hbar/\Omega$ for three different speeds marked by colored symbols in the main panel in (a): $\Delta_{\varepsilon}/\tau_{d}=(0.1,0.65,0.8,1.1,1.75)\Omega^{2}/\hbar$ and $\Delta\mu=0$. Units are set by the characteristic energy $\Omega$, which in quantum dot systems is of the order of meV.
@@
The second term, associated to the measurement [104, 105], appears as fluctuations of the first law [106] and vanishes upon averaging since $\hat{H}_{0},\hat{N}]=0$.
@@
Transport—The generated particle current, plotted in Fig. 2, is robust for a wide range of parameters, even against the bias when $\Delta\mu\leq 0$. In the adiabatic regime, where $\Delta_{\varepsilon}/\tau_{d}\to 0$, the charge transfer from l to r occurs with a high probability, $P_{LZ}\to 1$. Neglecting errors due to tunneling from R, the zero-bias current (so $f^{0}\equiv f^{0}=f_{R}^{0}$) can be estimated by $I(\Delta\mu=0)\approx[\tau_{d}+1/\Gamma f^{0}(1-f^{0})]^{-1}$ [107]. Hence the current is suppressed as $1/\tau_{d}$ for too slow drivings $\tau_{d}\gg\Gamma^{-1}$, see Fig. 2(a). As the system goes faster through the anticrossing, the current exhibits oscillations as a consequence of the coherent LZSM tunneling. The current increases, however it becomes less likely that the transition is done correctly in the nonadiabatic regime and the current saturates. This is reflected in the dispersion of the time averaged trajectories, ${I}_{\gamma}=\langle I^{\gamma}(t)\rangle_{t}$, plotted in Fig. 2(b): driving errors done by the demon make the current noisy [107]. In the extremely nonadiabatic regime, too many errors suppress the current [107].
@@
Conclusions—We have proposed a Maxwell demon that exploits the quantum properties of measurement and feedback in a simple transport setting. Undetailed detection of incoherently coupled subspaces (the total charge) enables a highly probable guess of the system state under appropriate preparation which allows a half-blind but confident feedback operation. The detector operator commutes with the system Hamiltonian, which averages the backaction out. Analyzing the quantum trajectories, we show that the properties of weakly nonadiabatic driving improve the transfer rate and the regularity of the generated current via operations where the demon performs no work on the system. The system entropy is hence reduced (with simultaneous cooling of the two reservoirs and the generation of electrical power) with the first law being respected, the only consumed resource being the continuous measurement.
@@
Acknowledgments—We thank Bibek Bhandari for useful comments on the manuscript. RS is grateful to Tobias Brandes for persistent inspiration. We acknowledge funding from the Spanish Ministerio de Ciencia e Innovación via grants No. PID2019-110125GB-I00 and No. PID2022-142911NB-I00, and through the “María de Maeztu” Programme for Units of Excellence in R&D CEX2023-001316-M.
@@
    where $\mathds{M}_{i}(\rho)$ is defined in Eq. (3). At each time step compute the survival probability
@@
    Immediately after detection, switch off tunneling ($\Gamma\to 0$) and sweep the dot energies linearly through the anticrossing. The state $\rho_{n}$ then evolves under
@@
We can write a simple model based on a rate equation for sequential transitions between an enlarged Hilbert space. It considers the empty state $\lvert 0\rangle$, and $\lvert l\rangle$, $\lvert r\rangle$ for states before the driving, and $\lvert l^{*}\rangle$, $\lvert r^{*}\rangle$ for states after the driving. The successful transitions $\lvert l\rangle\to\lvert r^{*}\rangle$ and $\lvert r\rangle\to\lvert l^{*}\rangle$ have a rate $\gamma_{LZ}=P_{LZ}/\tau_{d}$, and the errors $\lvert l\rangle\to\lvert l^{*}\rangle$ and $\lvert r\rangle\to\lvert r^{*}\rangle$ a rate $\bar{\gamma}_{LZ}=(1-P_{LZ})/\tau_{d}$. Assuming that the detection of tunneling through the contacts to the reservoirs is instantaneous, the transitions $\lvert L\rangle\to\lvert 0\rangle$ and $\lvert R\rangle\to\lvert 0\rangle$ are avoided, hence explicitly breaking detailed balance. We furthermore assume that the splitting is large enough for the interdot overlap to be negligible during the detection stages, $\Delta_{\varepsilon}\ll\lambda$ [118]. Then a local description in terms of localized states is convenient [103].
FIG. S1: Occupation probability of the right dot after going through the avoided crossing at a speed $\Delta_{\varepsilon}/\tau_{d}$, for $\lambda=0.5\Omega$, $\varepsilon_{0}=-0.5\Omega$ and $\Delta_{\varepsilon}=6\Omega$, for three different temperatures of the reservoirs, $k_{\text{B}}T=(10^{-3},0.75,2)\Omega$, and $\Delta\mu=0$. The lower temperature curves are almost identical. The inset zooms in the adiabatic regime.
@@
FIG. S2: Average particle current through the DQD computed with the heuristic model as a function of the applied electrochemical potential bias and the speed of the energy ramp, with cuts along the marked dotted lines plotted in the lateral panels. $\Gamma=\Omega/\hbar$, $\lambda=0.5\Omega$, $T_{l}=T_{r}=0.75\Omega/k_{\rm B}$, $\varepsilon_{0}=-0.5\Omega$, $\mu_{r}=0$, $\Delta=6\Omega$. In the lateral panels, the model is compared to the numerical results in Fig. 2 of the main text (grey lines). The black dotted line in the rightmost panel plots Eq. (S3).

### v2

# Noninvasive and nonadiabatic quantum Maxwell demon

@@ Abstract
Introduction—After decades of resting as a conceptual problem understood in terms of information erasure [1, 2], the interest on the Maxwell demon —a mechanism able to decrease the entropy of a system without acting directly on its particles [3]— has been revived in the last years. The way has been paved by the development of nanoscale technologies, with their access to microscopic degrees of freedom, bringing proposals and experimental realizations where the action of Maxwell demon devices [4, 5, 6, 7, 8] can actually be measured. This is so that a rich demonology has come out in the last years based on different platforms (e.g., quantum dots, ions, qubits), classified depending on whether they need active measurement and feedback on the system [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] or they work autonomously by means of mutual information flows in nonequilibrium configurations [19, 20, 21, 22, 23]. Information heat engines have also been proposed [24, 25, 26, 27, 28] and realized [29, 30] which are based on the exchange of mutual information [31] but cannot avoid the flow of heat.
@@
FIG. 1: Three steps of the demon operation inducing transport through a DQD coupled to reservoirs L and R: detection (of whether there is a particle in the DQD), operation (perform a voltage pulse that exchanges the energies of the left and right dot levels), and reset. The demon measures the total charge, NN, continuously and activates feedback operations when NN changes. $\Omega$ is the avoided crossing splitting.
@@
In the quantum regime, qubit based proposals [32, 33, 34, 35, 36, 37, 38, 39, 40] and proof of concept experiments [41, 42, 43, 44, 45] use quantum measurements [46, 47] and unitary gates. They involve either local projective measurements (so decoherence and work injection are unavoidable) or that the measurement and the feedback rotation are performed in different bases of the same qubit (so they cannot coexist). Here we solve these problems by proposing a Maxwell demon that exploits the quantum nature of the two demonic actions (detection and feedback) combining noninvasive continuous measurements and efficient driven tunneling [48] in a transport setting. It augments the qubit Hilbert space, ${\bf 1}=\{|l\rangle,|r\rangle\}$, by an empty state $|0\rangle$ that couples it to the work (or heat) extraction reservoirs and allows for continuous detection without perturbing the qubit.
@@
We exemplify this with a double quantum dot (DQD) charge qubit [49, 50] with interdot coupling $\Omega/2$ mediating the particle flow between two reservoirs (L and R) and coupled to a charge detector (e.g., a quantum point contact [51, 52, 53, 54], a single electron transistor [55] or a resonator [56, 57, 58, 59, 60]), see Fig. 1. The two states in $\bf 1$} are initially energetically split: $\varepsilon_{l}=\varepsilon_{0}$ and $\varepsilon_{r}=\varepsilon_{0}+\Delta_{\varepsilon}$, so the occupation of $\lvert l\rangle$ is favored. Upon the detection of an electron (most likely coming from L) the demon exchanges the level energies in a time $\tau_{d}$ (assumed to be much shorter than the characteristic dissipation time scale). During the level crossing, the electron tunnels to the other quantum dot with a Landau-Zener-Stückelberg-Majorana (LZSM) probability $P_{LZ}$ that is controlled by the driving speed, $\Delta_{\varepsilon}/\tau_{d}$ [61, 62, 63, 64]. By swapping both the dot energies and the charge configuration, the operation conserves energy and thereby enables transport: the electron can tunnel out to $R$. The initial state is then reset.
Provided $\Delta_{\varepsilon}\gg\Omega$, efficient $|l\rangle\rightarrow|r\rangle$ transfer is possible in the adiabatic ($\tau_{d}\gg\hbar/\Omega$) regime [65] as well as for particular speeds [66, 67]. LZSM tunneling is routinely used in DQDs to enable quantum gates [68, 69, 70, 71, 72, 73, 74, 75], and to probe coherence [76, 77, 78, 79, 80, 81, 82], also in qubits [83, 84, 85, 86]. Note that charge pumps generate transport using either periodic drivings [87, 88, 89, 90] without measuring, or conditioned feedback [91], in both cases with the drive performing work on the system [92]. To avoid this, the information of when the DQD gets occupied is essential.
@@
The challenge is to measure continuous and noninvasively in a way that the state coherence is not disturbed during the driving operation. Protocols conditioned on resolved knowledge of which dot ($l$ or $r$) is occupied [13, 16] impede the coherent driving and are affected by heat injection [93, 94, 95]. We avoid these drawbacks by imposing that the demon has only a limited access to the state of the system: it detects whether the DQD is empty ($N=0$) or occupied ($N=1$), but not in which quantum dot the electron is. Hence, using the detection operator $\hat{N}_{1}=\lvert l\rangle\langle l\rvert{+}\lvert r\rangle\langle r\rvert$ (the detector is equally coupled to both dots), the coherences between states $|l\rangle$ and $|r\rangle$ are respected during the operation step, even if the detector continues measuring. Note that the total charge $N=\langle\hat{N}\rangle$ plays the role of the memory of the demon in a Szilard engine [96, *szilard_on_1964]. By measuring $N=1$, the demon will hence not be certain that the electron is in the left quantum dot, as desired, but can prepare the system such that this is most likely the case if $\varepsilon_{0}\ll\varepsilon_{0}+\Delta_{\varepsilon}$.
@@
There is a price for indeterminacy: if the driving goes wrong (with probability $1-P_{LZ}$) the demon performs work and heats the system, which limits its performance far from the adiabatic regime; electrons tunneling from R will be transported by the demon in the wrong direction.
@@
$\hat{H}_{\mathrm{0}}(t)=\sum_{\alpha=l,r}\varepsilon_{\alpha}(t)\hat{n}_{\alpha}+\frac{\Omega}{2}\left(\hat{d}_{\mathrm{l}}^{\dagger}\hat{d}_{\mathrm{r}}+\hat{d}_{\mathrm{r}}^{\dagger}\hat{d}_{\mathrm{l}}\right).$ (1)
The time dependence is due to the demon action, according to the following protocol (see Fig. 1): (i) if $N=0$, $\varepsilon_{l}=\varepsilon_{0}$ and $\varepsilon_{r}=\varepsilon_{0}+\Delta_{\varepsilon}$ (detection I stage); (ii) by detecting a change to $N=1$ at $t=t_{k}$, the demon swaps the energies of the dots via gate voltages in a time $\tau_{d}$: $\varepsilon_{l}(t)=\varepsilon_{0}+\Delta_{\epsilon}(t-t_{k})/\tau_{s}$, $\varepsilon_{r}(t)=\varepsilon_{0}+\Delta_{\epsilon}(\tau_{d}t-t_{k})/\tau_{d}$ (driving stage); (iii) at $t=t_{k}+\tau_{d}$, $\varepsilon_{l}=\varepsilon_{0}+\Delta_{\varepsilon}$ and $\varepsilon_{r}=\varepsilon_{0}$ is maintained until $N=0$ is detected (detection II stage); (iv) finally a fast drive (compared to $1/\Gamma$) is performed to restore the initial state (reset stage). If $\varepsilon_{0}<\mu_{L}$ and $\mu_{R},k_{\text{B}}T\ll\varepsilon_{0}+\Delta_{\varepsilon}$, where $\mu_{j}$ is the electrochemical potential of reservoir $j=L,R$, tunneling into the DQD from R will be exponentially suppressed by a rate $\propto e^{-(\varepsilon_{0}+\Delta_{\varepsilon}-\mu_{R})/k_{\text{B}}T}$, so the demon can assume that most electrons come from L. To avoid further sources of errors, the demon also decouples the system from the reservoirs during the driving step, which can be done with additional gate voltages. We parametrize this as $\Gamma(t)=\Gamma\chi(t)$, where $\chi=0$ during the driving stage and 1 otherwise. In a successful sequence, the electron tunnels from L to $l$ at $t=t_{k}$, is transferred to $r$ at $t=t_{k}+\tau_{d}$ and then tunnels out to R. In that case, a particle is transferred across the system, heat $\mu_{j}-\varepsilon_{0}$ is extracted from both reservoirs and the demon has performed no work on the system. All the thermodynamic flows can be reversed in the system. The resource is in this case the amount of information generated in the detection apparatus to maintain the continuous measurement [98, 99, 100, 31].
We assume a fast and errorless detector. The evolution of the conditioned reduced density matrix of the DQD, $\rho_{\gamma}$ [101], for a finite time trajectory based on a sequence of measurements records ($\gamma_{(0,\tau)}$) is given by the stochastic master equation [102, 47]:
@@
where $L_{j}^{+}=\sqrt{\Gamma f_{j}}\lvert\alpha_{j}\rangle\langle 0\rvert$ and $L_{j}^{-}=\sqrt{\Gamma(1{-}f_{j})}\lvert 0\rangle\langle\alpha_{j}\rvert$ are jump operators associated with the tunneling in ($i=+$) and out ($i=$-) of the DQD via reservoir $j$, and $f_{j}=[e^{(\varepsilon_{\alpha_{j}}-\mu_{j})/k_{\text{B}}T}+1]^{-1}$ is the Fermi-Dirac distribution ($\alpha_{L}=l$, $\alpha_{R}=r$), which takes two values: $f_{j}^{0}$ and $f_{j}^{\Delta}$ depending on whether $\varepsilon_{\alpha_{j}}=\varepsilon_{0}$ or $\varepsilon_{0}+\Delta_{\varepsilon}$. Following Eq. (4), an initially empty DQD is updated to a state $\rho_{\gamma}^{\prime}=(f_{L}\lvert l\rangle\langle l\rvert+f_{R}\lvert r\rangle\langle r\rvert)/(f_{L}+f_{R})$. In the case $\Delta_{\varepsilon}\gg\mu_{R}+k_{\text{B}}T$ and $\mu_{L}-\varepsilon_{0}\sim k_{\text{B}}T$, the ratio $f_{L}/f_{R}\gg 1$ guarantees that the protocol will start with a state $\rho_{\gamma}^{\prime}\approx\lvert L\rangle\langle L\rvert$, with exponentially suppressed deviations, and $1-f_{j}$ is finite to enable the electron tunneling out to R after the driving. Additionally, the condition $\Delta_{\varepsilon}\ll\varepsilon_{0}$ suppresses the overlap of the quantum dot states when the system is not being driven, allowing for a local master equation description [103].
FIG. 2: (a) Average particle current through the DQD as a function of the applied electrochemical potential bias and the speed of the energy ramp, with cuts along the marked dotted lines plotted in the lateral panels, with $\Gamma=\Omega/\hbar$, $T_{l}=T_{r}=0.75\Omega/k_{\rm B}$, $\varepsilon_{0}=-0.5\Omega$, $\mu_{r}=0$, $\Delta=6\Omega$. The dashed line in the $\Delta\mu=0$ panel plots $\tau_{d}+1/\Gamma f^{0}(1-f^{0})]^{-1}$. (b) Histograms of the time-averaged current $I_{\gamma}\equiv\langle I^{\gamma}(t)\rangle_{t}$ of $10^{6}$ trajectories computed during a time $t_{run}=10^{5}\hbar/\Omega$ for three different speeds marked by colored symbols in the main panel in (a): $\Delta_{\varepsilon}/\tau_{d}=(0.1,0.65,0.8,1.1,1.75)\Omega^{2}/\hbar$ and $\Delta\mu=0$. We fix the splitting at the avoided crossing, $\Omega\sim 1\,\mathrm{meV}$
@@
The second term contains the dissipative contributions and the effect of the measurement [104, 102, 105], appears as fluctuations of the first law [106] and vanishes upon averaging since $\hat{H}_{0},\hat{N}]=0$.
@@
Transport—The generated particle current, plotted in Fig. 2, is robust for a wide range of parameters, even against the bias when $\Delta\mu\leq 0$. In the adiabatic regime, where $\Delta_{\varepsilon}/\tau_{d}\to 0$, the charge transfer from l to r occurs with a high probability, $P_{LZ}\to 1$. Neglecting errors due to tunneling from R, the zero-bias current (so $f^{0}\equiv f_{L}^{0}=f_{R}^{0}$) can be estimated by $I(\Delta\mu=0)\approx[\tau_{d}+1/\Gamma f^{0}(1-f^{0})]^{-1}$ [107]. Hence the current is suppressed as $1/\tau_{d}$ for too slow drivings $\tau_{d}\gg\Gamma^{-1}$, see Fig. 2(a). As the system goes faster through the anticrossing, the current exhibits oscillations as a consequence of the coherent LZSM tunneling. The current increases, however it becomes less likely that the transition is done correctly in the nonadiabatic regime and the current saturates. This is reflected in the dispersion of the time averaged trajectories, ${I}_{\gamma}=\langle I^{\gamma}(t)\rangle_{t}$, plotted in Fig. 2(b): driving errors done by the demon make the current noisy [107]. In the extremely nonadiabatic regime, too many errors suppress the current [107].
@@
Conclusions—We have proposed a Maxwell demon that exploits the quantum properties of measurement and feedback in a simple transport setting to ensure the system coherence and energy conservation. Undetailed detection of incoherently coupled subspaces (the total charge) enables a highly probable guess of the system state under appropriate preparation which allows for a half-blind but confident feedback operation. The detector operator commutes with the system Hamiltonian, which averages the backaction out. Analyzing the quantum trajectories, we show that the properties of weakly nonadiabatic driving improve the transfer rate and the regularity of the generated current via operations where the demon induces transport performing no work on the system. The system entropy is hence reduced (with simultaneous cooling of the two reservoirs and the generation of electrical power) with the first law being respected, the only consumed resource being the continuous measurement. Compared to classical analogues, the generated current is enhanced while noise is reduced.
@@
Acknowledgments—We thank Bibek Bhandari and Gonzalo Manzano for useful comments on the manuscript. RS is grateful to Tobias Brandes for persistent inspiration. We acknowledge funding from the Spanish Ministerio de Ciencia e Innovación via grants No. PID2022-142911NB-I00 and No. PID2024-157821NB-I00, and through the “María de Maeztu” Programme for Units of Excellence in R&D CEX2023-001316-M. L.T. acknowledges financial support from CSIC “JAE PRE” program (No. JAEPR23015), the CoQuSy project (No. PID2022-140506NB-C21 and C22), and the María de Maeztu Grant (No. CEX2021-001164-M) funded by MCIU/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR.
@@
    where $\mathds{M}_{i}(\rho)$ is defined in Eq. (3) of the main text. At each time step compute the survival probability
@@
    Immediately after detection, switch off tunneling ($\Gamma\to 0$) and sweep the dot energies linearly through the avoided crossing. The state $\rho_{n}$ then evolves under
@@
We can write a simple model based on a rate equation for sequential transitions between an enlarged Hilbert space. It considers the empty state $\lvert 0\rangle$, and $\lvert l\rangle$, $\lvert r\rangle$ for states before the driving, and $\lvert l^{*}\rangle$, $\lvert r^{*}\rangle$ for states after the driving. The successful transitions $\lvert l\rangle\to\lvert r^{*}\rangle$ and $\lvert r\rangle\to\lvert l^{*}\rangle$ have a rate $\gamma_{LZ}=P_{LZ}/\tau_{d}$, and the errors $\lvert l\rangle\to\lvert l^{*}\rangle$ and $\lvert r\rangle\to\lvert r^{*}\rangle$ a rate $\bar{\gamma}_{LZ}=(1-P_{LZ})/\tau_{d}$. Assuming that the detection of tunneling through the contacts to the reservoirs is instantaneous, the transitions $\lvert L\rangle\to\lvert 0\rangle$ and $\lvert R\rangle\to\lvert 0\rangle$ are avoided, hence explicitly breaking detailed balance. We furthermore assume that the splitting is large enough for the interdot overlap to be negligible during the detection stages, $\Delta_{\varepsilon}\gg\Omega$ [118]. Then a local description in terms of localized states is convenient [103].
FIG. S1: Occupation probability of the right dot after going through the avoided crossing at a speed $\Delta_{\varepsilon}/\tau_{d}$, for $\varepsilon_{0}=-0.5\Omega$ and $\Delta_{\varepsilon}=6\Omega$, for three different temperatures of the reservoirs, $k_{\text{B}}T=(10^{-3},0.75,2)\Omega$, and $\Delta\mu=0$. The lower temperature curves are almost identical. The inset zooms in the adiabatic regime.
@@
FIG. S2: Average particle current through the DQD computed with the heuristic model as a function of the applied electrochemical potential bias and the speed of the energy ramp, with cuts along the marked dotted lines plotted in the lateral panels. $\Gamma=\Omega/\hbar$, $T_{l}=T_{r}=0.75\Omega/k_{\rm B}$, $\varepsilon_{0}=-0.5\Omega$, $\mu_{r}=0$, $\Delta=6\Omega$. In the lateral panels, the model is compared to the numerical results in Fig. 2 of the main text (grey lines). The black dotted line in the rightmost panel plots Eq. (S3).